# Cross-Domain Few-Shot Learning with Coalescent Projections and Latent Space Reservation

### v1

@@ Abstract
Despite the progress in Cross-Domain Few-Shot Learning (CD-FSL), a model pre-trained with DINO combined with a prototypical classifier outperforms the latest SOTA methods. A crucial limitation that needs to be overcome is that updating too many parameters of the transformers leads to overfitting due to the scarcity of labeled samples. To address this challenge, we propose a new concept, Coalescent Projection (CP), as an effective successor to soft prompts. Additionally, we propose a novel pseudo-class generation method combined with Self-Supervised Transformations (SSTs) that relies solely on the base domain to prepare the network for encountering unseen samples from different domains. The proposed method exhibits its effectiveness in comprehensive experiments on the extreme domain shift scenario of the BSCD-FSL benchmark. Our code is published at https://github.com/Naeem-Paeedeh/CPLSR.
@@ 1 Introduction
Few-Shot Learning (FSL) has emerged as a novel approach to addressing data scarcity. The introduction of the BSCD-FSL benchmark [8] has, however, shown that many proposed methods are ineffective in practice when tested on a significantly different domain than the one on which the network was trained. For instance, they observed that all meta-learning methods underperform simple fine-tuning, and in some cases, those methods may underachieve compared to randomly initialized networks.
@@
We compared the reproducible State-Of-The-Art (SOTA) inductive methods on the BSCD-FSL benchmark, using a ViT-S backbone [5] pre-trained with DINO without further training, and evaluated the model on target datasets only with prototypes [24]. This is the pre-trained model used in StyleAdv [7] and PMF [11], which are the current SOTA methods. The results are presented in Fig. 1. The results indicate that SOTA methods still fall short of DINO.
@@
Ensuring that global features and local crops are mapped to the same region of the latent space helps the network attend to semantic features instead of domain-specific shortcuts  [38]. That is why DINO is difficult to beat. To achieve improvement over DINO, one needs to prepare a model to encounter new, unseen concepts while preserving the domain-invariant and class-agnostic patterns it learned in the past with DINO.
@@
An advantage of transformers over Convolutional Neural Networks (CNNs) that can be easily perceived in Natural Language Processing (NLP) is that they allow text prompts to provide context, controlling and adapting the frozen network’s behavior in new conditions. Nevertheless, manually designing the text prompts is time-consuming and fragile, as networks are susceptible to rephrasing and word choices. Later studies suggest this process can be automated by utilizing learnable continuous vectors instead of fixed text tokens [32, 17, 16]. These vectors of plain/soft prompts can point to better locations in the representation space, rather than the limited number of locations available in text token embeddings, thereby capturing the nuances of new, specific circumstances. Moreover, they can also be applied to other modalities, such as vision in Vision Transformers (ViTs) [5, 14], and can be inserted at any layer of a transformer, not just the first layer.
It is confirmed that using learnable plain prompts for every layer is even more effective than using them only for the first layer [1]. This finding suggests that the network’s attention to the input tokens should be redirected and corrected again with additional plain prompts at every layer. Therefore, the other computations in each block, such as the Multi-Layer Perceptron (MLP)’s computations, might be redundant, as we should rectify the main tokens again at the subsequent layers. Consequently, we aim to simplify what those prompts are intended to affect: the attention calculations and the input tokens.
To address the overfitting issue, we introduce CP as a successor and an improved alternative to plain prompts, offering several benefits. Since plain prompts can have an arbitrary length and only one or a few samples are available for each class in the target domain, finding the optimum number of prompt tokens is almost impossible, and one should rely only on heuristics or guessing. The first advantage of CP is that it does not introduce additional tokens. Second, it consumes less memory during training and inference, and it is more efficient. It is no longer required to calculate the interactions of extra tokens with other tokens in intermediate calculations. Next, CPs can control attention heads separately, which prevents their interference. Finally, it can be easily applied not only to the original transformer architectures and ViTs, but also to Swin Transformers [18] with minimal effort.
FAPLDP-NetPMFStyleAdvDINO40404040505050506060606070707070Avg. Acc.
@@
Figure 1: Comparison of SOTA inductive methods with DINO on BSCD-FSL benchmark. The blue and red bars indicate the 1-shot and 5-shot accuracies, respectively.
@@
To prepare the network to encounter samples from significantly different domains, we consider improving the network’s mapping in the latent space. Embeddings are semantically rich. Working with embeddings is very efficient, as we can perform calculations on the compressed, lower-dimensional vectors rather than the original inputs. Furthermore, we can treat the embeddings as equivalent to the original input samples to train the Neural Networks (NNs). In [21], it is shown that the text can be reconstructed from embeddings with high accuracy. Therefore, the embeddings can convey the same information, and one can augment the embeddings in such a way that the equivalent operation on the input space may not be easily explainable. Additionally, augmentations in the latent space do not require handcrafted transforms and are applicable to any modality.
@@
We propose pseudo-class generation to assist the network by compacting the area occupied by the base classes and reserving areas in the latent space for samples from the new domain, thereby separating them more easily and effectively. Specifically, we perform two types of augmentations in the base domain: the embedding level and the input level. At the embedding level, we generate pseudo-classes that enhance the network’s mapping by reserving parts of the latent space to anticipate unseen classes, while restricting the boundaries of the base classes. They enhance the mapping to create more complex and improved decision boundaries, thereby preventing the network’s reliance on patterns close to the base classes. At the input level, we deploy SSTs to generate additional novel classes and negative samples close to the current embeddings, thereby providing an additional repulsive force.
@@
    We propose the CP as a successor to learnable plain prompts. CPs achieve higher accuracies with lower memory requirements, control separate attention heads, do not require extra tokens, and can be applied to the Swin Transformers with negligible effort.
@@
    We propose a pseudo-class generation process that enables the network to anticipate unseen novel classes under domain shifts. That is, the pseudo-class generation mechanism guides in reserving representation spaces, thus adapting to an unseen domain seamlessly.
@@ 2 Related Works
Inductive methods in FSL, access only the support set and predict each query sample independently, which is a very challenging task, especially in CD-FSL.
@@
Wave-SAN [6] and FAP [35] focused on augmentation and utilizing the frequency domain. They separated the high and low-frequency components with the wavelet transform. In Wave-SAN, style augmentations were performed by swapping the low-frequency components by assuming they control the style and shape. In contrast, in FAP, the network’s reliance on high-frequency components was attempted to be reduced.
@@
The proposed method in [39] also operates in the frequency domain, but it utilizes a consistency constraint. StyleAdv [7] is an improvement over the Wave-SAN, and addresses the CD-FSL as a robustness issue. Instead of relying on the simple ”easy” styles generated by Wave-SAN, they adversarially generated the ”hard” to learn styles.
@@
Chen et al. [3] proposed an intra-block fusion to boost the extracted features in all convolution blocks in CNNs or Swin-S, and a cross-scale attention module to alleviate the scale-related inconsistencies due to the scarcity of the training data.
@@
PMF [11] demonstrated that utilizing modern architectures, such as ViTs, and self-supervised pre-training in combination with ProtoNet and fine-tuning can achieve very competitive performance. SemFew [34] uses semantics by exploiting the extra data from text modality to obtain more robust prototypes by aligning the vision and text embeddings of a Vision-Language Models (VLM).
@@
Transductive approach emerged under more relaxed conditions than the inductive approach. These methods can leverage both support and unlabeled query set samples during the inference [41, 37]. For instance, they can extract additional useful statistics, exploit similarities, distances, and structures to assign labels jointly and refine them to boost the performance. As a result, this approach is generally easier than the inductive approach.
@@
In APPL [9], the authors propose training a small parametric network that learns from the concatenation of the features of the support set samples for each class to generate prototypes, rather than relying on the average of embeddings as prototypes. protoLP [41] uses a novel prototype-based label propagation method and graph construction by considering the relation between the samples and prototypes instead of the relation between pairs of samples.
@@
IM-DCL [33] learns from the query set with a transductive mechanism and makes use of a distance-aware contrastive learning for a soft separation of the positive and negative sets. Dara [36] focuses on fast adaptation instead of domain generalization. It utilizes the query set’s statistics in a normalized distribution alignment module to solve the covariant shifts among the support and query samples.
@@
Self-Supvervised Learning (SSL) shows its effectiveness in transductive methods as additional unlabeled samples are available. The SWP [13] is proposed as a network pruning-based method that exploits the moderate number of unlabeled samples from the target domain through self-supervised classification. ADAPTER [22] pre-trains the model and aligns the domains with a bi-directional transformer architecture and DINO. Moreover, it uses label smoothing.
@@
In this paper, we focus on the inductive approach. As shown in Fig. 1, the SOTA methods are not better than DINO. Our analysis suggests that this may be due to the overfitting problem, as numerous parameters are involved in the training process. This background motivates us to develop a new method for the CD-FSL problem, which can be categorized as an inductive method.
@@ 3 Problem Formulation
The objective of CD-FSL is to classify samples belonging to unseen classes in the target domain $D_{T}$. Each task in CD-FSL is formulated as an $N$-way $K$-shot episode. In each episode, a support set $S=\{(x_{i},y_{i})\}_{i=1}^{NK}$ is created by drawing samples from $D_{T}$, where $x_{i}$ and $y_{i}$ are the $i$-th sample and its label, respectively, and $K$ is the number of samples per class. To evaluate the accuracy of a method, a query set $Q=\{(x_{i},y_{i})\}_{i=1}^{NM}$, drawn from $D_{T}$, is provided, where $M$ is the number of samples per query class. Since only a few samples from $N$ classes are insufficient for precisely estimating the accuracy of a method, we perform evaluation using a large number of episodes of such support and query sets to estimate the average accuracy.
@@ 4 The proposed method
Our method has two aspects. First, we introduce CP to rectify the attention of every head in the attention modules at every layer, which is more robust than plain prompts to overfitting. Second, we train those parameters to reserve the latent space by utilizing the pseudo-novel classes generated at both the input and latent space levels.
@@ 4.1 Coalescent Projection
Figure 2: The calculation of the plain prompts in the attention module on top and CP Prompts at the bottom. The blue color represents the input tokens, including the CLS token, and the yellow and gray elements indicate the learnable prompt tokens.
Let’s first examine what happens in the calculations of the plain prompts. Fig. 2 shows the intermediate calculations of both plain prompts and CP prompts. The elements of Region 1 are the result of the multiplication of query and key elements. The green Region 2 is controlled by prompts and query elements, and Region 3 is controlled by prompts and key elements. The elements of Region 4 are only being affected by prompt vectors. From AttnScale in [1], we know that scaling the attention map in Region 1 is a crucial part of the calculations in the attention. Moreover, the attention map is also where the domain alignment happens in cross-attention [22]. While the prompt tokens affect the area selected with red borders, including Regions 1 and 2, the calculations of the attention map can only be influenced by the image tokens in the query vector and plain prompts because of the softmax function.
@@
$\operatorname{Attn}(X)=\operatorname{Softmax}\left(\frac{\left(XW_{q}\right)\left(XW_{k}\right)^{T}}{\sqrt{d_{k}}}\right)XW_{v},$ (1)
@@
where $X$ denotes an $n\times d$ matrix of input tokens, $d$ denotes the dimension of features, $n$ stands for the number of tokens, $W_{q}$, $W_{k}$, and $W_{v}$ represent $d\times d_{k}$ matrices of query, key, and value projection weights, and $d_{k}$ is the attention heads’ dimension.
One straightforward way to adapt the critical part of attention (Region 1 in Fig. 2) without affecting the other parts is to add two separate projections and train those learnable parameters. However, it introduces too many parameters that may lead to overfitting. Instead, we propose utilizing a single weight matrix rather than having two separate matrices of weight vectors with additional parameters.
We assume that a hypothetical identity matrix $I\in\mathbb{R}^{d\times d}$ exists between the query and key matrices that does not alter the calculations as follows:
@@
$\operatorname{Attn}(X)=\operatorname{Softmax}\left(\frac{\left(XW_{q}\right)I\left(XW_{k}\right)^{T}}{\sqrt{d_{k}}}\right)XW_{v},$ (2)
To be able to steer the attention of the network, we set the non-diagonal elements to small non-zero values from a random normal distribution to create the CP matrix $C$ as follows:
$C_{ij}=\begin{cases}1,&\text{if}\quad i=j,\\ \varepsilon c_{ij},&\text{if}\quad i\neq j,\end{cases}$ (3)
where $c_{ij}\sim\mathcal{N}(0,\varepsilon)$. Finally, we obtain
$\operatorname{Attn}(X)=\operatorname{Softmax}\left(\frac{\left(XW_{q}\right)C\left(XW_{k}\right)^{T}}{\sqrt{d_{k}}}\right)XW_{v},$ (4)
We name it Coalescent Projection (CP)because it combines two concepts using a single projection matrix between them, rather than mapping the query and key intermediate embedding vectors with two projection matrices. We define a separate CP for each attention head to control the projected query and key vectors independently, thereby allowing each head to behave differently. The bottom graph in Fig. 2 shows the intermediate calculations for a single attention head with a CP.

### v2

@@ Abstract
Despite the progress in cross-domain few-shot learning, a model pre-trained with DINO combined with a prototypical classifier outperforms the latest SOTA methods. A crucial limitation that needs to be overcome is that updating too many parameters of the transformers leads to overfitting due to the scarcity of labeled samples. To address this challenge, we propose a new concept, coalescent projection, as an effective successor to soft prompts. Additionally, we propose a novel pseudo-class generation method, combined with self-supervised transformations, that relies solely on the base domain to prepare the network to encounter unseen samples from different domains. The proposed method exhibits its effectiveness in comprehensive experiments on the extreme domain-shift problem of the BSCD-FSL benchmark. Our code is published at https://github.com/Naeem-Paeedeh/CPLSR.
@@ 1 Introduction
Few-Shot Learning (FSL) has emerged as a novel approach to addressing data scarcity. However, the introduction of the challenging BSCD-FSL benchmark [guo2020broader] has shown that when one tests the FSL methods on a significantly different domain from that on which the network is trained, they become ineffective. For instance, it is observed that all meta-learning methods underperform simple fine-tuning and, in some cases, underachieve compared to randomly initialized networks [guo2020broader].
@@
We compared the reproducible State-Of-The-Art (SOTA) inductive methods on the BSCD-FSL benchmark, using a ViT-S backbone [dosovitskiy2020image], which was only pre-trained with DINO [caron2021emerging]. This pre-trained model was used in StyleAdv [fu2023styleadv] and PMF [hu2022pushing], the current SOTA methods. Next, we evaluated the model on target datasets with prototypes [Snell2017PrototypicalNF]. The results presented in Fig. 1 indicate that SOTA methods still fall short of DINO.
@@
DINO is difficult to beat because it ensures that the embeddings of global and local crops are mapped to the same region of the latent space, helping the network attend to semantic features rather than domain-specific shortcuts [zhou2023revisiting]. For further improvement, the model should be trained for new, unseen concepts while preserving the domain-invariant and class-agnostic patterns it learned with DINO.
@@
Transformers used for SOTA methods require more data than Convolutional Neural Networks (CNNs) [dosovitskiy2020image], but they possess powerful capabilities. In NLP, prepending prompts as instructions can provide context and control, and adapt the frozen network’s behavior to new conditions. However, networks are susceptible to rephrasing and word choices. Thus, this process is time-consuming and fragile. Later studies suggest that this process can be automated by utilizing learnable continuous vectors, named soft/plain prompts, instead of fixed text tokens [wang2022learning, li2021prefix, lester2021power]. These vectors can point to more precise locations in the latent space, rather than the limited number of locations available in text token embeddings, thereby capturing the nuances of new, specific circumstances. Moreover, they can also be applied to other modalities, such as vision in Vision Transformers (ViTs) [dosovitskiy2020image, jia2022visual], and can be inserted at any layer.
FAPLDP-NetPMFStyleAdvDINO4040505060607070Avg. Acc.1-shot5-shot
@@
Figure 1: Comparison of SOTA inductive methods with DINO on BSCD-FSL benchmark.
Using soft prompts for every layer is shown to be even more effective than restricting them to the first layer [basu2024strong]. This finding suggests that the network’s attention to the input tokens should be redirected and corrected again with additional soft prompts at every layer. We aim to simplify what those prompts are intended to affect: the attention map matrices and the input tokens.
We introduce Coalescent Projection (CP), an improvement over soft prompts that addresses overfitting and offers several benefits. For instance, finding the optimum number of prompt tokens is almost impossible, as only one or a few samples are available in FSL, and we should rely only on heuristics or guessing, but CP is independent of the number of tokens. As another example, CPs can handle attention heads separately, thereby preventing their interference. Moreover, it can change the relative ordering of original image tokens, while soft prompts cannot.
@@
To prepare a network to encounter unseen domains, we consider improving its embedding in the latent space. Embeddings are semantically rich, and working with them is very efficient, as we can perform calculations on the compressed, lower-dimensional vectors rather than the raw inputs. Furthermore, we can treat the embeddings as equivalent to the original input samples to train the Neural Networks (NNs). In [morris2023text], it is shown that the text can be reconstructed from embeddings with high accuracy. Therefore, the embeddings can convey the same information, and one can augment the embeddings in such a way that the equivalent operation on the input space may not be easily explainable. Additionally, augmentations in the latent space do not require handcrafted transforms and are applicable to any modality.
@@
We propose pseudo-class generation to assist the network by compacting the area occupied by base classes and reserving regions in the latent space for samples from new domains, thereby separating them more easily and effectively. Specifically, we perform two types of augmentations at the embedding and input levels. At the embedding level, we generate pseudo-classes that enhance the network’s mapping by reserving parts of the latent space to anticipate unseen classes, while restricting the boundaries of the base classes. They enhance mapping to create more complex and improved decision boundaries, thereby reducing the network’s reliance on patterns near the base classes. At the input level, we deploy Self-Supervised Transformations (SST) to generate additional novel classes and negative samples close to the current embeddings, thereby providing an additional repulsive force.
@@
    We propose the CP as a successor to soft prompts, achieving higher accuracies with low memory requirements, controlling separate attention heads, and without any extra tokens.
@@
    We propose a pseudo-class generation process that enables the network to anticipate unseen novel classes under domain shifts. That is, the pseudo-class generation mechanism guides the reservation of representation spaces, thus adapting to an unseen domain seamlessly.
@@ 2 Related Works
Inductive methods in FSL, access only the support set and predict each query sample independently, which is a very challenging task, especially in Cross-Domain Few-Shot Learning (CD-FSL).
@@
Wave-SAN [fu2022wave] and FAP [zhang2024exploring] focused on augmentation and utilizing the frequency domain. They separated the high and low-frequency components with the wavelet transform. In Wave-SAN, style augmentations were performed by swapping the low-frequency components, assuming they control the style and shape. In contrast, in FAP, the network’s reliance on high-frequency components was attempted to be reduced.
@@
The proposed method in [zhou2024meta] also operates in the frequency domain, but it utilizes a consistency constraint. StyleAdv [fu2023styleadv] is an improvement over the Wave-SAN, and addresses the CD-FSL as a robustness issue. Instead of relying on the simple ”easy” styles generated by Wave-SAN, they adversarially generated the ”hard” to learn styles.
@@
Chen et al[chen2024pushing] proposed an intra-block fusion to boost the extracted features in all convolution blocks in CNNs or Swin-S, and a cross-scale attention module to alleviate the scale-related inconsistencies due to the scarcity of the training data.
@@
PMF [hu2022pushing] demonstrated that utilizing modern architectures, such as ViTs, and self-supervised pre-training in combination with ProtoNet and fine-tuning can achieve very competitive performance. SemFew [zhang2024simple] uses semantics by exploiting the extra data from text modality to obtain more robust prototypes by aligning the vision and text embeddings of a Vision-Language Models (VLM).
@@
Transductive approach emerged under more relaxed conditions than the inductive approach. These methods can leverage both support and unlabeled query set samples during the inference [zhu2023transductive, Zhou2003LearningWL]. For instance, they can extract additional useful statistics, exploit similarities, distances, and structures to assign labels jointly and refine them to boost the performance. As a result, this approach is generally easier than the inductive approach.
@@
In APPL [heidari2024adaptive], the authors propose training a small parametric network that learns from the concatenation of the features of the support set samples for each class to generate prototypes, rather than relying on the average of embeddings as prototypes. protoLP [zhu2023transductive] uses a novel prototype-based label propagation method and graph construction by considering the relation between the samples and prototypes instead of the relation between pairs of samples.
@@
IM-DCL [xu2024enhancing] learns from the query set with a transductive mechanism and makes use of a distance-aware contrastive learning for a soft separation of the positive and negative sets. Dara [zhao2023dual] focuses on fast adaptation instead of domain generalization. It uses the query set’s statistics in a normalized distribution alignment module to address covariant shifts between the support and query samples.
@@
Self-Supvervised Learning (SSL) shows its effectiveness in transductive methods as additional unlabeled samples are available. The SWP [ji2024soft] is proposed as a network pruning-based method that exploits the moderate number of unlabeled samples from the target domain through self-supervised classification. ADAPTER [paeedeh2024cross] pre-trains the model and aligns the domains with a bi-directional transformer architecture and DINO, and uses label smoothing.
@@
In this paper, we focus on the inductive approach. As Fig. 1 shows, the SOTA methods are not better than DINO. Our analysis suggests this may be due to overfitting, given the numerous parameters involved in the training process. Moreover, the base samples are similar to those used in pre-training. This background motivates us to develop a new inductive method for the CD-FSL problem.
@@ 3 Problem Formulation
The objective of CD-FSL is to classify samples belonging to unseen classes in the target domain $D_{T}$ by having one or a few samples for each class. Each task in CD-FSL is formulated as an $N$-way $K$-shot episode. In each episode, a support set $S=\{(x_{i},y_{i})\}_{i=1}^{NK}$ is created for training by drawing samples from $D_{T}$, where $x_{i}$ and $y_{i}$ are the $i$-th sample and its label, respectively, and $K$ is the number of samples per class. To evaluate the accuracy of a method, a query set $Q=\{(x_{i},y_{i})\}_{i=1}^{NK_{q}}$, drawn from $D_{T}$, is provided, where $K_{q}$ is the number of samples per query class. Since only a few samples from $N$ classes are insufficient for precisely estimating the accuracy of a method, we perform evaluation using a large number of episodes of such support and query sets to estimate the average accuracy.
@@ 4 The proposed method
Our method has two major components. First, we introduce CP to calibrate each head’s attention in the attention module at each layer, making it more robust to overfitting than soft prompts. Second, we train those parameters to reserve the latent space by leveraging pseudo-novel classes generated at both the input and latent space levels.
@@ 4.1 Coalescent Projection
Figure 2: From top, the calculation of the soft prompts, AttnScale, and CP, in the attention module. $\odot$ is the Hadamard product.
In this section, we start by studying the intermediate calculations of the soft prompts in the attention mechanism. Next, after finding the crucial part of the calculations, we propose CP as a better alternative.
@@
$\operatorname{Attn}(X)=\operatorname{Softmax}\left(\frac{QK^{\text{T}}}{\sqrt{d_{k}}}\right)V,$ (1)
@@
where $Q=XW_{q}$, $K=XW_{k}$, and $V=XW_{v}$ represent the query, key, and value matrices after projections, and $d_{k}$ is the attention heads’ dimension.
Fig. 2 shows the intermediate calculations of soft prompts, AttnScale, and CP for a single head. The attention map matrix for soft prompts, depending on whether they are influenced by the input tokens or the prompt tokens, comprises four blocks:
$M=\frac{QK^{\text{T}}}{\sqrt{d_{k}}}=\begin{bmatrix}Q_{\text{I}}K_{\text{I}}^{\text{T}}&Q_{\text{I}}K_{\text{P}}^{\text{T}}\\ Q_{\text{P}}K_{\text{I}}^{\text{T}}&Q_{\text{P}}K_{\text{P}}^{\text{T}}\\ \end{bmatrix}=\\ \begin{bmatrix}M_{\text{I,I}}&M_{\text{I,P}}\\ M_{\text{P,I}}&M_{\text{P,P}}\\ \end{bmatrix}$. (2)
From $\tilde{M}$ calculations in AttnScale [basu2024strong], we know that scaling the elements of the $M_{\text{I,I}}$ block (block 1 in Fig. 2) in the attention map is the crucial part of the calculations. Moreover, the attention map is also where the domain alignment happens in cross-attention [paeedeh2024cross]. Furthermore, the prompt tokens can only influence the $M_{\text{I,I}}$ and $M_{\text{I,P}}$ (blocks 1 and 2 with red borders) because of the existence of the softmax function. Since new prompt tokens in subsequent layers should correct the effect of prompts from previous layers, we consider the bottom blocks redundant.
A more direct way to adapt and influence the $M_{\text{I,I}}$ is to train the query and key projections. However, two matrices have too many fine-tuning parameters, which leads to overfitting and forgetting. Instead, we propose using a single learnable matrix that can be viewed as a projection in both directions of the query and key sides, with other parameters kept frozen.
We consider a learnable square matrix $C$ between the query and key matrices:
@@
$\operatorname{Attn}(X)=\operatorname{Softmax}\left(\frac{QCK^{\text{T}}}{\sqrt{d_{k}}}\right)V,$ (3)
and name it Coalescent Projection (CP) because it combines and unites two concepts (projected query and key matrices here) to coalesce them into a single concept (attention map) using a single projection matrix, rather than mapping them with two projection matrices (query and key matrices). If $\Delta W_{q}$ and $\Delta W_{k}$ are the ideal changes to the query and key projections after training on the new domain, to achieve the same effect, $C$ can be set to:
$C=I+\Delta W_{k}^{\text{T}}{W_{k}^{\text{T}}}^{\dagger}+W_{q}^{\dagger}\Delta W_{q}+W_{q}^{\dagger}\Delta W_{q}\Delta W_{k}^{\text{T}}{W_{k}^{\text{T}}}^{\dagger},$ (4)
Please see the supplementary material for the proof. We utilize CP for each attention head, allowing the heads to be handled independently without interference.
4.1.1 CP vs. Soft Prompts
CP has crucial advantages over prompts. First, the number of prompt tokens is a crucial hyperparameter that cannot be precisely determined before comprehensive testing. Second, the prompt tokens are shared across all heads; their gradients can interfere with one another, which is detrimental to optimization and generalization. In contrast, CP can handle heads independently. Besides, the growth in the number of tokens across consecutive layers when using soft prompts increases the tokens’ interactions, further confusing the network. Next, adding new prompt tokens in each layer gradually consumes the context window and memory. Finally, CP can change the order of the original input tokens, whereas soft prompts cannot, as we define it next.
Axiom 1.
The prerequisite for changing the order of tokens is to be able to change two elements of row $i$ in the attention map $M$.
Theorem 1.
CP can change the relative ordering of original image tokens, whereas soft prompts cannot.
The proof is provided in the supplementary material.
Corollary 1.
The prompts can redistribute some probability mass to the prompt values $V_{ij}$, $n\leq i\leq n+p$, which are then propagated to later layers.
4.1.2 CP vs. AttnScale
AttnScale as a SOTA Parameter-Efficient Fine-Tuning (PEFT) method outperformed prevalent and effective PEFT methods, Adapter, LoRA, and prompts, on Meta-Dataset on [basu2024strong]. Here, we compare CP to AttnScale.
CP is compatible with any number of input tokens as it is always $d_{k}\times d_{k}$. By contrast, the size of the AttnScale matrix is $n\times n$ for $n$ input tokens, and it is not applicable for different token lengths, such as those resulting from adding prompts or when the model encounters variable-length sequences in NLP.
CP is more efficient when a higher-resolution image or longer sequence is given, and it is less prone to overfitting. For a network with $L$ layers and $n_{h}$ heads for each attention, the space complexity of CP is $O(Ln_{h}d_{k}^{2})$, whereas AttnScale requires $O(Ln_{h}n^{2})$. For example, in the ViT-S in our experiments, AttnScale requires $(\frac{n}{d_{k}})^{2}=(\frac{197}{64})^{2}\approx 9.5\times$ of the CP parameters.
Finally, CP can stretch or shrink in the latent space because CP has the ability to learn with the Mahalanobis distance metric, while AttnScale cannot (please refer to the supplementary material for details)